Lavage Rapide is a Canadian company that owns and operates a large automatic car wash facility near Montreal. The following table provides data concerning the company’s costs: Fixed Cost per Month Cost per Car Washed Cleaning supplies $ 0.70 Electricity $ 1,100 $ 0.07 Maintenance $ 0.20 Wages and salaries $ 4,500 $ 0.30 Depreciation $ 8,400 Rent $ 1,800 Administrative expenses $ 1,400 $ 0.01 For example, electricity costs are $1,100 per month plus $0.07 per car washed. The company expects to wash 8,000 cars in August and to collect an average of $6.80 per car washed. The company actually washed 8,100 cars in August.

Required: Calculate the company's activity variances for August. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)

Someone commented on this saying it needed more information but this is the entire question, please answer.
Answer #1

Answer:

Particulars Budgeted Actual Activity variance F/U
Cars 8000 8100
Sales $             54,400 $           55,080 $          680 F
Costs:
Cleaning supplies $               5,600 $             5,670 $            70 U
Electricity $               1,660 $             1,667 $               7 U
Maintenance $               1,600 $             1,620 $            20 U
Wages and salaries $               6,900 $             6,930 $            30 U
Depreciation $               8,400 $             8,400 $             -   None
Rent $               1,800 $             1,800 $             -   None
Administrative expenses $               1,480 $             1,481 $               1 U
Total costs $             27,440 $          27,568 $          128 U
Net Income $             26,960 $          27,512 $          552 F

Calculations:

- А в с DE 284 Particulars Budgeted Actual Activity variance F/U 285 286 Cars 8000 8100 287 Sales =6.8*B286 =6.8*0286 =C287-B

In case of any doubt, please feel free to comment.